# Testing the dark side of neutrino oscillations with the solar neutrino fog at Dark Matter experiments

### v1

@@ I Introduction
Recently, the first observations of CEvNS with solar neutrinos at dark matter (DM) direct detection experiments has been reported by the large-scale xenon experiments XENONnT and PandaX-4T with more than significance Aprile et al. (2024a); Bo et al. (2024). This so-called “neutrino fog” originates from 8B solar neutrinos and provides an irreducible background at dark matter direct detection experiments Monroe and Fisher (2007); Vergados and Ejiri (2008); Strigari (2009); Gutlein et al. (2010); Billard et al. (2014); O’Hare (2021); Akerib et al. (2022). In the Standard Model (SM), the CEvNS process is mediated by the boson however, in scenarios beyond the SM where neutrinos have non-standard interactions (NSI) new contributions to the CEvNS cross sections can arise. In Dutta et al. (2017); Bœhm et al. (2019); Schwemberger and Yu (2022); Amaral et al. (2023) the effect of new interactions on the neutrino fog has been studied and the PandaX-4T and XENONnT data sets have been analyzed by the neutrino community to probe new neutrino interactions in Aristizabal Sierra et al. (2024); Li et al. (2024); Blanco-Mas et al. (2024); Maity and Boehm (2024); Demirci and Mustamin (2025); De Romeri et al. (2024); Karadag et al. (2025).
@@
Apart from addressing the question of how “normal” neutrinos are and determining one of the last unknown parameters in the 3-flavor neutrino paradigm Denton et al. (2022), determining the neutrino mass ordering has important implications for other experiments, in particular for experiments which measure observables related to the absolute neutrino mass scale. For example, the predicted parameter space for the beta decay end point spectrum observable is lower for NO than for IO, affecting the reach of the KATRIN experiment Osipowicz et al. (2001) and future experiments Monreal and Formaggio (2009); Gastaldo et al. (2017); Alpert et al. (2015). Current cosmological observatories like DESI Aghamousa et al. (2016), the Atacama Cosmology Telescope Calabrese et al. (2025), and the South Pole Telescope Camphuis et al. (2025) are close to measuring the sum of the neutrino masses; in fact, the strongest bounds from DESI are in conflict with the minimal allowed value in IO Elbers et al. (2025). Unambiguously determining the neutrino mass ordering from oscillation experiments will be essential to determine if there is possible new physics in cosmology or even in oscillation physics. Also the detection prospects of the cosmic neutrino background by capture on tritium depends on the mass ordering Long et al. (2014). Finally, the allowed parameter space for neutrinoless double beta decay experiments allows much smaller minimal values of depending on the lightest mass in NO than in IO by at least one order of magnitude in the hierarchical region of neutrino masses Denton and Gehrlein (2024). Correctly identifying the mass ordering is therefore essential for neutrinoless double beta decay experiments in determining their required sensitivity.
In the following we will first review the LMA-Dark degeneracy and its current status in sec. II (see Denton and Gehrlein (2022) for a more detailed overview). In sec. III we discuss neutrino scattering at DM direct detection experiments before we outline our analysis in sec. IV. Our results and next steps in fully ruling out the LMA-Dark degeneracy are described in sec. V and we conclude in sec. VI.
@@ II LMA-Dark degeneracy
where is the neutrino energy, is the lepton mixing matrix Pontecorvo (1957); Maki et al. (1962), is related to the matter potential, is Fermi’s constant, and is the electron number density. In the presence of NSI all entries in the second matrix can be populated. The NSI parameters in the Hamiltonian are related to the NSI parameters in the Lagrangian from eq. (1) as
@@
Constraints on NSI from oscillations have been derived in Esteban et al. (2024) and no strong evidence for non-zero NSI has been found. However, the oscillation probability derived from eq. (2) suffer from a degeneracy in the presence of NSI, called the LMA-Dark degeneracy de Gouvea et al. (2000); Bakhti and Farzan (2014); Coloma and Schwetz (2016); Denton and Parke (2022). Namely, as a consequence of CPT invariance, the Hamiltonian transforms as , which corresponds to a change in the oscillation parameters as
@@
with for .222Depending on the definition of the mass eigenstates, the LMA-Dark degeneracy can also be defined as instead of Denton (2020); Denton and Parke (2022). If then the LMA-Dark degeneracy is fully contained in the electron neutrino sector, if the LMA-Dark degeneracy is in the muon and tau neutrino sectors. In turn, the LMA-Dark degeneracy also means that one cannot measure the mass ordering unambiguously with oscillation experiments. Here an alternative way to test NSI comes to the rescue, namely neutrino neutral current scattering experiments. These experiments do not suffer from the LMA-Dark degeneracy but have their own degeneracies, unrelated to LMA-Dark. Since matter effects in oscillations are coherent forward scattering processes, they are independent of the mediator mass of the new interaction. In contrast, in scattering processes the mediator mass plays a role. Therefore, to probe the LMA-Dark degeneracy one has to test all mediators down to light masses. Indeed, cosmology excluded neutrinophilic mediators with mass below the MeV scale. For a vector mediator the bound reads MeV Kamada and Yu (2015); Huang et al. (2018); Blinov et al. (2019); Sabti et al. (2021); Li and Xu (2023). So far, the LMA-Dark region has been excluded for a mediator with masses above the GeV scale with couplings to quarks and electrons using NuTeV and CHARM data where neutral current deep inelastic scattering is dominant Coloma et al. (2017a, b). At lower energies the dominant neutral current process is CEvNS. Using CEvNS data sets from COHERENT, the LMA-Dark region with mediator masses above 50 MeV and coupling to quarks and electron or muon neutrinos have been excluded in Denton et al. (2018); Chaves and Schwetz (2021). For lighter mediators, CEvNS data from the reactor experiment Dresden-II has excluded the LMA-Dark region for couplings between electron neutrinos and quarks Denton and Gehrlein (2022). Additionally recent data from CONUS+ Ackermann et al. (2024); De Romeri et al. (2025); Chattaraj et al. (2025a) excluded this parameter space as well, see sec. V. The LMA-Dark parameter space for the coupling of electron neutrinos to electrons down to mediator masses constrained by cosmology has been excluded by solar neutrino experiments like Borexino Dutta et al. (2020); Agostini et al. (2018), and reactor experiments like TEXONO, CHARM-II, and GEMMA Lindner et al. (2018); Deniz et al. (2010); Vilain et al. (1994); Beda et al. (2010). As there are currently no constraints on the LMA-Dark parameter space for the coupling of muon and tau neutrinos to electrons or quarks, this remains the only viable region for LMA-Dark.333Very specific combinations of couplings of electron neutrinos to up and down quarks are allowed as well Denton and Gehrlein (2022).
@@
There are two possible ways to probe LMA-Dark in the muon and tau neutrino sector: Low threshold experiments using neutrinos from pion decay at rest sources like Coherent CAPTAIN Mills Aguilar-Arevalo et al. (2022), or at the ESS Baxter et al. (2020); Chattaraj et al. (2025b) or SNS Adamski et al. (2024, 2025) can constrain directly the coupling of a new mediator to muon neutrinos and quarks Shoemaker and Welch (2021); Chaves and Schwetz (2021) while the muon neutrino-electron elastic scattering cross section is several orders of magnitude smaller. Combined with information from oscillation experiments which allow only for a small difference in the new interactions of muons and tau neutrinos with quarks Coloma et al. (2023) constraints on new interactions in the tau sector can be derived. Here we will follow an alternative route which allows to directly probe new interactions of muon and tau neutrinos simultaneously, namely the observation of solar neutrinos in a low energy scattering processes. Since solar neutrinos arrive at the Earth dominantly as mass eigenstate they contain a large fraction of the muon neutrino and tau neutrino flavor eigenstates. A study of the couplings of muon and tau neutrinos to electrons is possible using solar neutrino-electron scattering data Coloma et al. (2022a); Brdar and Xu (2023); Kelly et al. (2024); Mishra and Strigari (2023) like current Borexino data Appel et al. (2022) and future JUNO data Abusleme et al. (2023) and DUNE data Capozzi et al. (2019) as well as neutrino-electron scattering in DM direct detection experiments. Here we will focus on testing the new interactions of muon and tau neutrinos with quarks using the CEvNS process instead.
@@
Note that SNO’s observation of solar neutrino NC scattering on deuterium Ahmad et al. (2002); Aharmim et al. (2013) is dominated by the axial-vector current whereas the LMA-Dark degeneracy appears for vector current only. Therefore we need to resort to low energy scattering processes with a vector current like CEvNS. In the next section we analyze the 8B solar neutrino scattering data from PandaX-4T and XENONnT and show that current data does not exclude LMA-Dark yet and outline possible future experimental configurations to test LMA-Dark.
@@ III Neutrino scattering at DM direct detection experiments
The dominant contribution to the solar neutrino fog are neutrinos originating from the process in the core of the Sun. Neutrinos from the process have similar energies but their flux is three orders of magnitude smaller than for 8B neutrinos Vinyoles et al. (2017) and hence do not contribute significantly to the neutrino fog. After their production, due to the high matter densities in the Sun and matter effects, the 8B neutrinos undergo an adiabatic flavor transition where they propagate as an asymptotic propagation eigenstate before they exit the Sun Wolfenstein (1978). Since this eigenstate contains all three neutrino flavor eigenstates, this means that 8B neutrinos provide a source of low energy muon and tau neutrinos without the need to rely on meson, muon, or tau decays.
In the presence of NSI, in general, the flavor transition probability in the Sun gets affected. However, the LMA-Dark parameter space in the Sun does not change the oscillation probability. Therefore, we do not include the effects of NSI in the propagation of the neutrinos through the Sun. In fig. 1 we show the neutrino survival and transition probabilities for 8B neutrinos using the PEANUTS code Gonzalo and Lucente (2024), along with the flux of 8B neutrinos. The shape of the 8B spectrum was obtained from Bahcall et al. (1996); Bahcall , and the normalization was obtained from the B16-GS98 Standard Solar Model Vinyoles et al. (2017).
@@
The diffraction radius is calculated using a folding width fm Lewin and Smith (1996) and the rms charge radius of xenon taken from Angeli and Marinova (2013). Finally, is the weak charge in the presence of NSI. For the case of only diagonal non-zero NSI, it is given by
@@
For light mediators whose mass is below the momentum transfer of the scattering MeV) eq. (10) we have to replace444We note that in some references a factor of 2 appears in the denominator of this expression, however as demonstrated in the appendix of Atzori Corona et al. (2022) this factor is erroneous.
@@ IV Analysis and Results
Turning now to our analysis of the data from XENONnT Aprile et al. (2024b) and PandaX-4T Meng et al. (2021). Both experiments are dual-phase time projection chambers featuring 5.9 t and 3.7 t of liquid xenon as targets. XENONnT across its two runs SR0 and SR1, with a live time of 108 days and 208.5 days respectively, reported the observation of 37 total events with a background expectation of 26.4, leading to an observation of 8B solar neutrinos at 2.73. The experiment has a threshold of 0.5 keV and a 3.51 t-yr exposure across both runs. PandaX-4T also presented results from two runs, the commissioning run, and the first science run of the experiment. PandaX-4T did two separate analysis for their 8B CEvNS data, one where they required coincidence of S1 and S2 signals (paired), and another where they used ionization signals only (US2). We only use the US2 data in this analysis due to the higher statistics. PandaX-4T reports observing 75 8B neutrino events in their US2 data with a total exposure of 1.04 t-yr. XENONnT presents the results in cS2 bins, which are the corrected number of photoelectrons (PE) from the ionization signal. The electron gain = 16.9 PE/electron Aprile et al. (2024a) for XENONnT relates the number of ionization electrons released to the number of PE observed in the cS2 signal. The event rate in each electron/cS2 bin is calculated using a Poisson smearing of events. The number of events in the bin is
@@
when is the mean number of electrons/PE released for an event with nuclear recoil energy for PandaX-4T/XENONnT respectively, calculated using the charge yields from Ma et al. (2023); Aprile et al. (2025). is the exposure of the respective experiment, with t-yr for XENONnT, and t-yr for PandaX-4T US2 dataset. is the efficiency of the experiments taken from Bo et al. (2024) for PandaX-4T and from Aprile et al. (2024a) for XENONnT. We use a total of three bins for XENONnT of size (90,160], (160,230], (230,500] PE and eight bins of equal width between [4,8) electrons for PandaX-4T. We find an excellent agreement for the SM event rates for PandaX-4T and a slightly worse but still acceptable agreement for XENONnT which may be due to details in the experimental analysis.

### v2

@@ I Introduction
Recently, the first observations of CEvNS with solar neutrinos at dark matter (DM) direct detection experiments has been reported by the large-scale xenon experiments XENONnT and PandaX-4T with more than significance Aprile et al. (2024a); Bo et al. (2024). This so-called “neutrino fog” originates from 8B solar neutrinos and provides an irreducible background at dark matter direct detection experiments Monroe and Fisher (2007); Vergados and Ejiri (2008); Strigari (2009); Gutlein et al. (2010); Billard et al. (2014); O’Hare (2021); Akerib et al. (2022). In the Standard Model (SM), the CEvNS process is mediated by the boson however, in scenarios beyond the SM where neutrinos have nonstandard interactions (NSI) new contributions to the CEvNS cross sections can arise. In Dutta et al. (2017); Bœhm et al. (2019); Schwemberger and Yu (2022); Amaral et al. (2023) the effect of new interactions on the neutrino fog has been studied and the PandaX-4T and XENONnT data sets have been analyzed by the neutrino community to probe new neutrino interactions in Aristizabal Sierra et al. (2024); Li et al. (2024); Blanco-Mas et al. (2024); Maity and Boehm (2024); Demirci and Mustamin (2025); De Romeri et al. (2024a); Karadag et al. (2025); De Romeri et al. (2024b, c).
@@
Apart from addressing the question of how “normal” neutrinos are and determining one of the last unknown parameters in the three-flavor neutrino paradigm Denton et al. (2022), determining the neutrino mass ordering has important implications for other experiments, in particular for experiments which measure observables related to the absolute neutrino mass scale. For example, the predicted parameter space for the beta decay end point spectrum observable is lower for NO than for IO, affecting the reach of the KATRIN experiment Osipowicz et al. (2001) and future experiments Monreal and Formaggio (2009); Gastaldo et al. (2017); Alpert et al. (2015). Current cosmological observatories like DESI Aghamousa et al. (2016), the Atacama Cosmology Telescope Calabrese et al. (2025), and the South Pole Telescope Camphuis et al. (2025) are close to measuring the sum of the neutrino masses; in fact, the strongest bounds from DESI are in conflict with the minimal allowed value in IO Elbers et al. (2025). Unambiguously determining the neutrino mass ordering from oscillation experiments will be essential to determine if there is possible new physics in cosmology or even in oscillation physics. Also the detection prospects of the cosmic neutrino background by capture on tritium depends on the mass ordering Long et al. (2014). Finally, the allowed parameter space for neutrinoless double beta decay experiments allows much smaller minimal values of depending on the lightest mass in NO than in IO by at least one order of magnitude in the hierarchical region of neutrino masses Denton and Gehrlein (2024). Correctly identifying the mass ordering is therefore essential for neutrinoless double beta decay experiments in determining their required sensitivity.
In the following we first review the LMA-Dark degeneracy and its current status in sec. II (see Denton and Gehrlein (2022) for a more detailed overview). In sec. III we discuss neutrino scattering at DM direct detection experiments before we outline our analysis in sec. IV. Our results and next steps in fully ruling out the LMA-Dark degeneracy are described in sec. V and we conclude in sec. VI.
@@ II LMA-Dark degeneracy
where is the neutrino energy, is the PMNS mixing matrix Pontecorvo (1957); Maki et al. (1962) that is parameterized in the usual way Navas et al. (2024), is related to the Wolfenstein matter potential Wolfenstein (1978), is Fermi’s constant, and is the electron number density. In the presence of NSI all entries in the second matrix can be populated. The NSI parameters in the Hamiltonian are related to the NSI parameters in the Lagrangian from eq. (1) as
@@
Constraints on NSI from oscillations have been derived in Esteban et al. (2024) and no strong evidence for nonzero NSI has been found. However, the oscillation probability derived from eq. (2) suffers from a degeneracy in the presence of NSI, called the LMA-Dark degeneracy de Gouvea et al. (2000); Bakhti and Farzan (2014); Coloma and Schwetz (2016); Denton and Parke (2022). Namely, as a consequence of CPT invariance, the Hamiltonian transforms as , which corresponds to a change in the oscillation parameters as
@@
with for .222Depending on the definition of the mass eigenstates, the LMA-Dark degeneracy can also be defined as instead of Denton (2020); Denton and Parke (2022). If then the LMA-Dark degeneracy is fully contained in the electron neutrino sector, if the LMA-Dark degeneracy is in the muon and tau neutrino sector.
@@
In turn, the LMA-Dark degeneracy also means that one cannot measure the mass ordering unambiguously with oscillation experiments. Here an alternative way to test NSI comes to the rescue, namely neutrino neutral current scattering experiments. These experiments do not suffer from the LMA-Dark degeneracy but have their own degeneracies, unrelated to LMA-Dark. Since matter effects in oscillations are coherent forward scattering processes, they are independent of the mediator mass of the new interaction. In contrast, in scattering processes the mediator mass plays a role. Therefore, to probe the LMA-Dark degeneracy one has to test all mediators down to light masses. Indeed, cosmology excluded neutrinophilic mediators with mass below the MeV scale. For a vector mediator the bound reads MeV, depending on the data sets combined Kamada and Yu (2015); Huang et al. (2018); Blinov et al. (2019); Sabti et al. (2021); Li and Xu (2023); Ghosh et al. (2024).333Our model also predicts neutrino self-interactions due to the new mediator which can be constrained using supernova data Fiorillo et al. (2023, 2024a, 2024b); Akita et al. (2022); Telalovic et al. (2024). So far, the LMA-Dark region has been excluded for a mediator with masses above the GeV scale with couplings to quarks and electrons using NuTeV and CHARM data where neutral current deep inelastic scattering is dominant Coloma et al. (2017a, b). At lower energies the dominant neutral current process is CEvNS. Using CEvNS data sets from COHERENT, the LMA-Dark region with mediator masses above 50 MeV and coupling to quarks and electron or muon neutrinos has been excluded in Denton et al. (2018); Chaves and Schwetz (2021). For lighter mediators, CEvNS data from the reactor experiment Dresden-II has excluded the LMA-Dark region for couplings between electron neutrinos and quarks Denton and Gehrlein (2022). Additionally recent data from CONUS+ Ackermann et al. (2024); De Romeri et al. (2025); Chattaraj et al. (2025a) excluded this parameter space as well, see sec. V. The LMA-Dark parameter space for the coupling of electron neutrinos to electrons down to mediator masses constrained by cosmology has been excluded by solar neutrino experiments like Borexino Dutta et al. (2020); Agostini et al. (2018), and reactor experiments like TEXONO, CHARM-II, and GEMMA Lindner et al. (2018); Deniz et al. (2010); Vilain et al. (1994); Beda et al. (2010). As there are currently no constraints on the LMA-Dark parameter space for the coupling of muon and tau neutrinos to electrons or quarks, this remains the only viable region for LMA-Dark.444Very specific combinations of couplings of electron neutrinos to up and down quarks are allowed as well Denton and Gehrlein (2022).
@@
There are two possible ways to probe LMA-Dark in the muon and tau neutrino sector: Low threshold experiments using neutrinos from pion decay at rest sources like Coherent CAPTAIN Mills Aguilar-Arevalo et al. (2022), or at the ESS Baxter et al. (2020); Chattaraj et al. (2025b) or SNS Adamski et al. (2024, 2025) can constrain directly the coupling of a new mediator to muon neutrinos and quarks Shoemaker and Welch (2021); Chaves and Schwetz (2021) while the muon neutrino-electron elastic scattering cross section is several orders of magnitude smaller. Combined with information from oscillation experiments which allow only for a small difference in the new interactions of muons and tau neutrinos with quarks Coloma et al. (2023) constraints on new interactions in the tau sector can be derived. Here we follow an alternative route which allows to directly probe new interactions of muon and tau neutrinos simultaneously, namely the observation of solar neutrinos in a low energy scattering processes. Since solar neutrinos arrive at the Earth dominantly as mass eigenstate they contain a large fraction of the muon neutrino and tau neutrino flavor eigenstates Wolfenstein (1978). A study of the couplings of muon and tau neutrinos to electrons is possible using solar neutrino-electron scattering data Coloma et al. (2022a); Brdar and Xu (2023); Kelly et al. (2024); Mishra and Strigari (2023) like current Borexino data Appel et al. (2022) and future JUNO data Abusleme et al. (2023) and DUNE data Capozzi et al. (2019) as well as neutrino-electron scattering in DM direct detection experiments. Here we will focus on testing the new interactions of muon and tau neutrinos with quarks using the CEvNS process instead.
@@
Note that SNO’s observation of solar neutrino neutral current scattering on deuterium Ahmad et al. (2002); Aharmim et al. (2013) is dominated by the axial-vector current whereas the LMA-Dark degeneracy appears for the vector current only. Therefore we need to resort to low energy scattering processes with a vector current like CEvNS. In the next section we analyze the 8B solar neutrino scattering data from PandaX-4T and XENONnT and show that current data does not exclude LMA-Dark yet and outline possible future experimental configurations to test LMA-Dark.
@@ III Neutrino scattering at DM direct detection experiments
The dominant contribution to the solar neutrino fog is neutrinos originating from the process in the core of the Sun. Neutrinos from the process have similar energies but their flux is three orders of magnitude smaller than for 8B neutrinos Vinyoles et al. (2017) and hence do not contribute significantly to the neutrino fog. After their production, due to the high matter densities in the Sun and matter effects, the 8B neutrinos undergo an adiabatic flavor transition where they propagate as an asymptotic propagation eigenstate before they exit the Sun, as shown in the fundamental paper by Wolfenstein Wolfenstein (1978). Since this eigenstate contains all three neutrino flavor eigenstates, this means that 8B neutrinos provide a source of low energy muon and tau neutrinos without the need to rely on meson, muon, or tau decays.
In the presence of NSI, in general, the flavor transition probability in the Sun gets affected. However, the LMA-Dark parameter space in the Sun does not change the oscillation probability. Therefore, we do not include the effects of NSI in the propagation of the neutrinos through the Sun. In fig. 1 we show the neutrino survival and transition probabilities for 8B neutrinos using the PEANUTS code Gonzalo and Lucente (2024), along with the flux of 8B neutrinos. The shape of the 8B spectrum was obtained from Bahcall et al. (1996); Bahcall , and the normalization was obtained from the B16-GS98 Standard Solar Model Vinyoles et al. (2017). While the survival probability for electron neutrinos is accurately known, the individual transition probabilities and depend sensitively on and which are currently the least well measured parameters of the leptonic mixing matrix Esteban et al. (2024). For concreteness we use the following values for the parameters of the PMNS matrix , , , and . For this set of parameters the transition probabilities and are identical.
@@
The diffraction radius is calculated using a folding width fm Lewin and Smith (1996) and the rms charge radius of xenon taken from Angeli and Marinova (2013). Finally, is the weak charge in the presence of NSI. For the case of only diagonal nonzero NSI, it is given by
@@
For light mediators whose mass is below the momentum transfer of the scattering MeV) eq. (10) we have to replace555We note that in some references a factor of 2 appears in the denominator of this expression, however as demonstrated in the appendix of Atzori Corona et al. (2022) this factor is erroneous.
@@ IV Analysis and Results
We turn now to our analysis of the data from XENONnT Aprile et al. (2024b) and PandaX-4T Meng et al. (2021). Both experiments are dual-phase time projection chambers featuring 5.9 t and 3.7 t of liquid xenon as targets. XENONnT across its two runs SR0 and SR1, with a live time of 108 days and 208.5 days respectively, reported the observation of 37 total events with a background expectation of 26.4, leading to an observation of 8B solar neutrinos at 2.73. The experiment has a threshold of 0.5 keV and a 3.51 t-yr exposure across both runs. PandaX-4T also presented results from two runs, the commissioning run, and the first science run of the experiment. PandaX-4T did two separate analyses for their 8B CEvNS data, one where they required coincidence of S1 and S2 signals (paired), and another where they used ionization signals only (US2). We only use the US2 data in this analysis due to the higher statistics. PandaX-4T reports observing 75 8B neutrino events in their US2 data with a total exposure of 1.04 t-yr.
XENONnT presents the results in cS2 bins, which are the corrected number of photoelectrons (PE) from the ionization signal. The electron gain = 16.9 PE/electron Aprile et al. (2024a) for XENONnT relates the number of ionization electrons released to the number of PE observed in the cS2 signal. The event rate in each electron/cS2 bin is calculated using a Poisson smearing of events. The number of events in the bin is
@@
when is the mean number of electrons/PE released for an event with nuclear recoil energy for PandaX-4T/XENONnT respectively, calculated using the charge yields from Ma et al. (2023); Aprile et al. (2025). is the exposure of the respective experiment, with t-yr for XENONnT, and t-yr for PandaX-4T US2. is the efficiency of the experiments taken from Bo et al. (2024) for PandaX-4T and from Aprile et al. (2024a) for XENONnT. We use a total of three bins for XENONnT of size (90,160], (160,230], (230,500] PE and eight bins of equal width between [4,8) electrons for PandaX-4T. We find an excellent agreement for the SM event rates for PandaX-4T and a slightly worse but still acceptable agreement for XENONnT which may be due to details in the experimental analysis.